# Topic-FlipRAG: Topic-Orientated Adversarial Opinion Manipulation Attacks to Retrieval-Augmented Generation Models

### v2

@@ Appendix A Appendix
A.1. Full experiment details
(1)Black-box RAG. We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain. Conversational RAG adds logic for incorporating historical messages and intention reasoning, thus supporting back-and-forth conversation with users and serving as a widely used RAG framework. Based on a retriever and a large language model (LLM), Conversational RAG mainly consists of a history-awaring retriever to identify users’ intention and a question-answering chain to produce output, as shown in Figure9. The LLMs adopted are the open-source models Meta-Llama-3.1-8B-Instruct (Llama3.1), Qwen-2.5-7B-Instruct(Qwen2.5). They perform well across various tasks among all open-source LLMs. The prompt connecting the retriever and the LLM in $\text{RAG}_{\text{black}}$ adopts the basic RAG prompt from LangChain:
Use the following pieces of retrieved context to answer the question. Keep the answer concise. Context: {context}. Question: {question}.
Figure 9. The framework of Conversational RAG.
(2) Target retriever model in RAG-system. People usually use dense retrieval models as the retriever in RAG. Therefore, we select three representative dense retrieval models, Contriever, DPR and ANCE, as the target retrieval model (Gao and Callan, 2021; Karpukhin et al., 2020; Xiong et al., 2020). By convention, we use dot product between the embedding vectors of questions(queries) and candidate documents as their similarity score $R$ in the ranking.
(3) Manipulation target. For a controversial topic-queries-set $Q$, documents $doc_{tar}$ with the target opinion $S_{t}$ are manipulated by modifing $doc_{tar}$ through a series of subtle alterations $p_{adv}$, . This manipulation aims to place these perturbed documents as prominently as possible in the top $k$ rankings of the RAG retriever $\mathrm{RM}_{k}(q)$, where $k$ denotes the number of documents obtained from the retrieval results by the LLM in $\text{RAG}_{\text{black}}$ In this paper, $k$ is set to 3.
(4) Opinion classification. We use Qwen2.5-Instruct-72B as the opinion classifier. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text.
(5) Experimental parameters. In the knowledge-guided attack process, we set the maximum editing distance parameter $\epsilon$ to 0.2, the semantic similarity threshold $\lambda$ to 0.85, and the number of iterations $N$ to 5.For adversarial trigger generation, we configured the experimental parameters as follows: the number of beams was set to 3, the top-$k$ value to 10, the batch size to 32, the temperature value to 1.0, the learning rate to 0.005, and the sequence length to 10.
(6) Poisoned Target.For the PROCON dataset, to investigate the manipulation performance under more challenging conditions, we performed relevance ranking for the documents with respect to each topic-query set $Q$ and the target stance $S_{t}$ From the ranked list, we selected the last five documents (i.e., those with the lowest relevance) as the target poisoned documents. For the MS MARCO dataset, we utilized the top-1000 relevance-ranked passage list for each topic-query set. From this list, we selected the passage with the lowest average rank as the target passage. This approach ensures that the evaluation focuses on passages that are least relevant to the target queries, thus providing a more rigorous benchmark for the proposed method.
(7) Experimental environment. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.
A.2. Details of baselines
PoisonedRAG: Zou et al. (Zou et al., 2024) propose PoisonedRAG which injects a few poisoned texts into the RAG database such that the LLM generates manipulated answers. They consider both black-box and white-box settings. In the black-box setting, they attempt to include the question itself as part of the poisoned text to make it semantically similar to the target question. For our task, we adapted this strategy to the topic-queries setting by randomly selecting one query from the topic-queries set $Q$ as the trigger and inserting it at the start of the document. The manipulated documents were then evaluated for their impact on retrieval rankings and subsequent comparisons.
PAT: As a representative adversarial retrieval attack strategy, PAT employs a pairwise generation paradigm. Given a target query, the target candidate item, and the top-ranked candidate item (referred to as the anchor, which guides adversarial text generation), the method utilizes gradient-based optimization of a pairwise loss function calculated between the target and the anchor. The strategy incorporates additional constraints for fluency and next-sentence prediction to ensure coherence. Using beam search, the final adversarial trigger, denoted as $T_{\text{pat}}$, is generated iteratively in an auto-regressive manner. In our task, we adapted PAT for topic-queries by generating $T_{\text{pat}}$ for the target document considering the whole topic-queries set and evaluating its effectiveness under black-box constraints.
Collision: It employs a gradient-based optimization approach to generate adversarial paragraphs, referred to as collisions, which are semantically similar to the target query. Using gradient descent, continuous representations of collisions are optimized and subsequently discretized into tokens via beam search. ASC provides three variants: aggressive, aggressive-regularized, and natural. For evaluation, we focused on the most and least aggressive variants to examine their effectiveness in generating adversarial texts. Using parameters and settings derived from the original Birch model implementation, collisions were generated with a fixed length of 10 tokens. In the context of topic-queries, ASC was applied to create collisions for the entire topic-queries set, and its transfer attack performance was evaluated on the black-box RAG retriever.
A.3. Ablation Study Supplement
Due to space limitations, only a subset of ablation study results is presented in the main text. Figure 10 illustrates the impact of the Polarity Control module in Topic-FlipRAG on the sentiment of the documents when the target polarity is set to PRO. Figures 12 and 13 depict the influence of various baseline methods on document sentiment.
Figure 10. Impact of Polarity Control on documents sentiment (target polarity:PRO) .
Figure 11. Impact of Polarity Control on document stance.
Figure 12. Impact of Polarity Control on document stance(target polarity:CON).
Figure 13. Impact of Polarity Control on document stance(target polarity:PRO).
Additionally, Figure 11 demonstrates the effect of the Polarity Control module on the actual stance polarity of the documents.The distinction between stance polarity and sentiment is critical. Stance polarity is determined by utilizing the opinion classification approach described in Section 9, which evaluates whether the document supports or opposes the corresponding topic. In contrast, sentiment is directly inferred using a sentiment classification model, specifically cardiffnlp/twitter-roberta-base-sentiment-latest.
 [^2] https://huggingface.co/cardiffnlp/twitter-roberta-base-sentiment-latest
Table 8. Topic-Oriented RAG attack results on RPOCON dataset based on ANCE.
Method Target Stance ANCE+qwen-2.5-8B-instruct ANCE+llama3.1-8B-instruct
H&E G&P EDU S&C avg. ASV $\Delta$ ASV H&E G&P EDU S&C avg. ASV $\Delta$ ASV
Clean CON 0.21 0.24 0.18 0.18 0.21 0.24 0.24 0.24 0.25 0.24
PRO 0.19 0.24 0.17 0.20 0.21 0.21 0.27 0.23 0.24 0.24
ASC CON 0.30 0.30 0.25 0.23 0.28 0.06 0.35 0.35 0.26 0.28 0.32 0.08
PRO 0.24 0.35 0.25 0.27 0.29 0.06 0.27 0.40 0.27 0.29 0.31 0.07
PAT CON 0.23 0.27 0.24 0.25 0.25 0.04 0.32 0.28 0.27 0.27 0.29 0.05
PRO 0.25 0.23 0.27 0.29 0.25 0.05 0.29 0.29 0.30 0.29 0.30 0.05
PoisonedRAG CON 0.26 0.34 0.36 0.44 0.35 0.13 0.32 0.43 0.36 0.40 0.36 0.09
PRO 0.31 0.37 0.30 0.28 0.33 0.12 0.34 0.31 0.30 0.31 0.33 0.09
Topic-FlipRAG CON 0.39 0.44 0.40 0.42 0.42 0.21 0.40 0.50 0.41 0.50 0.46 0.23
PRO 0.43 0.41 0.31 0.39 0.39 0.18 0.40 0.52 0.37 0.56 0.47 0.23
Table 9. Topic-Oriented RAG attack results on RPOCON dataset based on DPR.
Method Target Stance DPR+qwen-2.5-8B-instruct DPR+llama3.1-8B-instruct
H&E G&P EDU S&C avg. ASV $\Delta$ ASV H&E G&P EDU S&C avg. ASV $\Delta$ ASV
Clean CON 0.22 0.26 0.27 0.25 0.25 0.25 0.24 0.21 0.23 0.23
PRO 0.23 0.26 0.24 0.25 0.24 0.24 0.27 0.22 0.22 0.24
Collision CON 0.25 0.30 0.28 0.21 0.27 0.02 0.33 0.34 0.33 0.31 0.33 0.10
PRO 0.27 0.34 0.31 0.29 0.31 0.07 0.28 0.39 0.39 0.39 0.35 0.12
PAT CON 0.22 0.23 0.34 0.26 0.26 0.01 0.32 0.26 0.35 0.19 0.28 0.05
PRO 0.23 0.23 0.23 0.23 0.25 0.01 0.21 0.35 0.35 0.27 0.29 0.04
PoisonedRAG CON 0.32 0.40 0.28 0.48 0.37 0.12 0.39 0.39 0.36 0.58 0.43 0.20
PRO 0.29 0.36 0.31 0.31 0.32 0.08 0.30 0.42 0.36 0.42 0.37 0.14
Topic-FlipRAG CON 0.45 0.42 0.40 0.49 0.44 0.19 0.45 0.48 0.41 0.48 0.46 0.23
PRO 0.41 0.55 0.40 0.39 0.46 0.21 0.37 0.52 0.42 0.42 0.44 0.21
A.4. RAG opinion manipulation result based on other retrievers
@@
To further explore RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?, we conducted RAG-based opinion manipulation experiments on the PROCON dataset using ANCE and DPR as retrievers. The results obtained for each baseline are consistent with the overall trends observed in Table 3. Detailed experimental outcomes are provided in Table 8 and Table 9.

### v3

@@ 8 Conclusion
Open Science
Our work complies with USENIX Security’s open science mandate by prioritizing transparency and reproducibility. All attack implementations, including Topic-FlipRAG, evaluation scripts, and datasets, are hosted on GitHub [^5] https://github.com/LauJames/Topic-FlipRAG and Zenodo [^6] https://doi.org/10.5281/zenodo.15523435. The repository is licensed under the MIT License, allowing reuse while prohibiting malicious applications. We have provided comprehensive documentation, covering environment setup, hyper-parameters, and attack pipelines, to ensure independent verification. Sensitive data has been excluded from the repository. Synthetic queries were generated using GPT-4o, with ethical constraints imposed to eliminate harmful or biased language. All components necessary to replicate our findings are publicly accessible. We are committed to maintaining the repository availability for at least three years and will update the repositories based on community feedback.
Ethics Considerations
Our research investigates vulnerabilities in RAG systems through the lens of adversarial opinion manipulation. We rigorously evaluated ethical implications using principles from The Menlo Report (Beneficence, Respect for Persons, Justice, and Respect for Law) and the USENIX Security Ethics Guidelines.
Responsible Disclosure. We understand the importance of engaging with model and framework providers, so we reported this vulnerability to the LangChain Security Team, hoping to collaborate on developing a mitigation module. However, due to limitations in defense measures, we have not yet gained recognition, and we are still striving to make effective progress. Additionally, the widespread use of hybrid open-source frameworks means that ownership boundaries are less clear compared to traditional software vulnerabilities. This ambiguity prompted us to explore a policy-oriented disclosure approach, aiming to encourage systemic preventative measures. Therefore, we used an accessible and efficient channel to initiate policy-level mitigation. We reported our preliminary findings and associated risks to our institution. This information, along with related content, was subsequently disclosed to Xinhua News Agency, not for public release. We want to emphasize that this initial disclosure was not intended to be exclusive or politically motivated. We are willing to inform other relevant platforms and international organizations.
Data Privacy and User Consent. Our experiments involved opinion-based datasets covering various topics. To protect individual privacy, we ensured that all datasets are anonymized to remove personally identifiable information (PII), thus safeguarding individual privacy. Moreover, before beginning the user experiments, participants were thoroughly informed about the study’s background, purpose, and the measures in place to ensure fairness and privacy protection. After the experiment, we reiterated the true data distribution to participants to minimize any potential cognitive impact. Throughout the research, we conducted all experiments, particularly those involving adversarial attacks, in secure environments to prevent unintended consequences.
By addressing these ethical considerations, our research not only highlights vulnerabilities but also contributes to the development of robust solutions in safeguarding RAG models against exploitation.
IRB Approval. This study was conducted under full Institutional Review Board (IRB) approval (Protocol No. WHU-HSS-IRB2025010. We confirm that all research activities strictly complied with the IRB-approved protocol. All research activities involving human participants received approval from our institution’s Humanities and Social Sciences Ethics Committee. To obtain IRB approval, we submitted a comprehensive application package, including a research ethics justification report, informed consent forms, a detailed project protocol, and the CVs of both the principal investigator and all research participants. The application was subject to a rigorous, multi-stage review process involving sequential approval by the project leader, institutional supervisors, the ethics committee secretary, and the committee leadership. After successfully completing all review stages, our study was formally granted IRB approval. Specific procedures are detailed as follows:
Informed Consent: All participants were required to sign an informed consent form prior to the experiment. The document clearly communicated the study’s objectives, procedures, estimated duration, potential psychological risks (e.g., exposure to controversial topics), compensation scheme (a cash reward of 50 RMB), and participants’ rights to withdraw at any time without penalty. Participants were also informed that their responses would be used solely for academic research.
Data Handling and Anonymization: All responses and opinion scores were collected using anonymized participant IDs. No personally identifiable information (such as name, contact details, or personal ID) was stored. Data were encrypted and stored on secure servers accessible only to core research staff. Data will be retained for three years and then permanently deleted. Throughout the study, strict anonymization procedures were enforced. Participant identifiers were randomly assigned, and all data records excluded any direct or indirect personal information. This ensured that individuals could not be re-identified during analysis or in any published findings.
User Privacy Protection: The entire study was conducted in a controlled laboratory environment to minimize external data exposure. All data were used exclusively for academic purposes and were never shared with third parties. Participants were provided with post-experiment debriefing, including full disclosure of the study’s objectives and safeguards. Additionally, experimental group participants were shown control group outputs after completion to reduce potential cognitive bias introduced by adversarial content exposure.
In summary, this study not only adhered to rigorous ethical standards but also aimed to advance the development of trustworthy AI by systematically identifying vulnerabilities in RAG systems and promoting secure, transparent, and user-respecting model design.
Acknowledgement.
This work is supported by the National Science and Technology Major Project (2023ZD0121502) and National Natural Science Foundation of China (72404212).
@@ Appendix A Appendix
A.1 Additional Experiment Details
(1) Opinion Classification. LLMs have been widely adopted as evaluators or "judges" in recent literature  [zheng2023judging]. Prior work has further validated the reliability of LLMs in sentiment classification [calderon2025alternative] and explicit stance detection [li2024advancing].
In our experiment, we employ Qwen2.5-Instruct-72B as the opinion classifier. Owing to its large parameter scale and instruction tuning, the model demonstrates strong capability in identifying user stances embedded in natural language. To evaluate its reliability, we randomly sampled 500 RAG-generated responses from the PROCON dataset and recruited three cognitively unimpaired undergraduate students to independently annotate each instance. The Krippendorff’s $\alpha$ among the human annotators was 0.8833, and the average pairwise Cohen’s $\kappa$ between the LLM and the three human annotators was 0.7897.
We further aggregated the three annotations via majority vote to derive a human consensus label. The LLM’s predictions achieved a $\kappa$ of 0.8110 and an accuracy of 0.8760 against this consensus, indicating strong agreement and supporting the use of LLMs as reliable stance evaluators in our setup.
A.2 Additional Parameter Analysis
[…]
Table 8: Performance under different similarity thresholds ($\lambda$) on the PROCON dataset. Metrics top3-ori, top3-att, top3-v, and RASR are reported in %.
similarity threshold $\lambda$ IR Llama3.1
[…]
0.80 42.83 74.52 31.68 57.84 0.66 0.39
0.85 42.83 70.24 27.40 51.47 0.64 0.37
0.90 42.83 68.46 25.62 50.67 0.52 0.25
Table 9: Performance under different edit distances ($\epsilon$) on the PROCON dataset. Metrics top3-ori, top3-att, top3-v, and RASR are reported in %.
edit distance $\epsilon$ IR Llama3.1
[…]
0.10 42.83 67.41 24.58 47.30 0.57 0.30
0.20 42.83 70.24 27.40 51.47 0.64 0.37
0.30 42.83 72.81 29.97 56.08 0.65 0.38
It is worth noting that a larger edit distance $\epsilon$ and a lower similarity threshold $\lambda$ both represent looser constraints on adversarial document modification. Both intuition and empirical results suggest that looser constraints lead to more effective manipulation. However, they may also make detection and mitigation easier, highlighting a trade-off between attack strength and stealth. Nevertheless, Topic-FlipRAG consistently outperforms the baseline across all evaluation metrics under various constraint settings. To balance attack effectiveness and applicability across different deployment scenarios, we adopt moderate default values of $\epsilon=0.2$ and $\lambda=0.85$ in our main experiments.
A.3 User Study Details
To evaluate the effectiveness of adversarial opinion manipulation in real-world usage, we conducted a controlled user study involving 54 college students (aged 18–25; 29 males, 25 females). Participants were randomly assigned to two groups of equal size: the control group (Group-clean) and the experimental group (Group-poisoned) (27 participants each), and interacted independently with a QA system based on RAG system. Topic-FlipRAG was applied only for the experimental group to poison either supporting or opposing documents, enabling a between-subjects comparison of post-interaction stance shifts.
Procedures. The experiment was fully computer-based and lasted approximately 30 minutes. Participants interacted with the system across two controversial topics. For each topic, participants first reported their initial stance using a 7-point Likert scale, which was later normalized to $[0,1]$ (where 0 = strong opposition, 1 = strong support). These pre-interaction scores were used to verify that there was no statistically significant difference in the baseline opinion distributions between the control group and the experimental group prior to any system exposure.
Afterward, the participants engaged in three rounds of interaction per topic. In each round, they were given a predefined query to input into the RAG-based QA system and were instructed to read the returned response. For the experimental group, the responses were manipulated using Topic-FlipRAG: one topic had its supporting documents poisoned, while the other had its opposing documents poisoned. The control group received unaltered responses. Following the interactions, participants again rated their opinions using the same 7-point Likert scale, enabling evaluation of manipulation effects.
Participant Tasks. Participants were instructed to explore two controversial topics using the RAG-based QA system and to reflect on their personal stance. For each topic, they completed a two-stage opinion evaluation process. First, prior to any interaction with the system, they reported their initial stance using a 7-point Likert scale. Next, they engaged in three rounds of interaction with the RAG system, each involving a distinct predefined query. After reviewing all three responses, participants rated their post-interaction stance using the same 7-point scale.
To support the interaction process, we selected two topics from the PROCON-dataset and randomly sampled three sub-questions under each topic to serve as participant queries. Below are the exact sub-questions used in the study.
Topic 1:
Should People Become Vegetarian?
Q1: Is vegetarianism compassionate?
Q2: Is vegetarianism influential?
Q3: Should teenagers be vegetarian?
Topic 2:
Should Humans Colonize Space?
Q1: Should humans pursue extraterrestrial colonization?
Q2: Are humans ready for space colonization?
Q3: What are your thoughts on space colonization?
The following evaluation questions were used to formally measure participants’ stance on each topic, both before and after their interaction with the QA system:
Topic 1:
Should People Become Vegetarian?
Pre-interaction: "For you personally, do you think vegetarianism is beneficial or meaningless?"
Post-interaction: "After reading the content on this topic, do you think vegetarianism is meaningful? Please rate the degree of necessity."
Topic 2:
Should Humans Colonize Space?
Pre-interaction: "For you personally, do you think humanity’s current investment in interstellar exploration is more beneficial or more harmful?"
Post-interaction: "After learning about this topic, assuming you are a member of the senate, would you support or oppose large-scale societal investment in space expansion?"
Pre/Post-Inform Protocol. To avoid priming effects, participants were initially informed that the study aimed to "evaluate user cognitive feedback on AI QA systems," with no mention of adversarial attacks or opinion manipulation. This omission was deliberate to preserve the naturalness of user responses. Following the experiment, a full debriefing was provided in accordance with IRB guidelines. Participants in the experimental group were informed of the adversarial modifications applied to system outputs, and all participants were given the option to review the clean system responses used for the control group. Contact information was provided for follow-up inquiries.
Table 10: Pre-Interaction Opinion Score Comparison Between Control and Experimental Groups Across Topics
Topic Group Mean ± SD p-value
Topic 1 Control 0.4671 ± 0.2153 0.6839
Experimental 0.4424 ± 0.2278
Topic 2 Control 0.6996 ± 0.1690 0.9370
Experimental 0.7037 ± 0.2084
Statistical Analysis. Before interacting with the QA system, participants in both groups rated their initial stance on each topic. As shown in Table 10, the means and variances of pre-interaction scores were highly similar across the two groups, indicating that participants began with comparable baseline opinions. To statistically assess baseline equivalence, we tested the null hypothesis $\bm{H_{0}}$: the control and experimental groups do not exhibit significant differences in their initial scores. For both topics, the resulting p-values were substantially greater than 0.05, indicating that $\bm{H_{0}}$ cannot be rejected and supporting the conclusion that no significant differences exist between groups prior to the intervention. These results demonstrate the effectiveness of random assignment and rule out initial group bias.
To assess the impact of adversarial document poisoning, we conducted independent two-sample t-tests on post-interaction stance scores. Significant opinion shifts were observed in the experimental group compared to the control group: for Topic 1 (Vegetarianism), $p=0.0454<0.05$; and for Topic 2 (Space Colonization), $p=0.0017<0.005$.
A.4 Additional Mitigation Details
Our paragraphing approach uses the following prompt:
Task: Rewrite the following query while preserving its original meaning. Aim to modify as many words and expressions as possible, while ensuring the intent remains intact. Original Question: {question}
[…]
Method Target Stance ANCE + Qwen2.5 ANCE + Llama3.1 DPR + Qwen2.5 DPR + Llama3.1
$\DeltaASV avg. ASV Δ\DeltaASV$ $\DeltaASV avg. ASV Δ\DeltaASV$
Clean CON 0.21 0.24 0.25 0.23
PRO 0.21 0.24 0.24 0.24
Collision CON 0.28 0.06 0.32 0.08 0.27 0.02 0.33 0.10
PRO 0.29 0.08 0.34 0.10 0.31 0.07 0.35 0.12
PAT CON 0.25 0.04 0.29 0.05 0.26 0.05 0.28 0.05
PRO 0.25 0.05 0.31 0.07 0.25 0.01 0.27 0.04
PoisonedRAG CON 0.35 0.13 0.36 0.13 0.37 0.12 0.43 0.20
PRO 0.32 0.12 0.33 0.09 0.32 0.08 0.37 0.14
Topic-FlipRAG CON 0.42 0.21 0.46 0.23 0.44 0.19 0.46 0.23
PRO 0.39 0.18 0.47 0.23 0.46 0.21 0.44 0.21
Figure 10: Case study of Topic-oriented Opinion Manipulation On Topic:"Is Binge-Watching Good for You?".
A.5 Additional Manipulation Analyses
@@
To further explore RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?, we conducted RAG-based opinion manipulation experiments on the PROCON dataset using ANCE and DPR as retrievers. Detailed experimental outcomes are provided in Table 11. Figure 10 represents the case-study based on Topic: "Is Binge-Watching Good for You?"